OpenStudy (anonymous):

"ella took her boyfriend to a pizza store and spent 24.50 on 2 buffets and 2 hours of gaming. the next night she took her parents, buying 3 buffets and 4 hours of gaming and spend 37. 25. what is the cost of a buffet meal?

OpenStudy (anonymous):

x = buffet y = gaming 2x + 2y = 24.50 || 4x + 4y = 49 3x + 4y = 37.25 || 3x + 4y = 37.25 x = 11.75

OpenStudy (anonymous):

how did you do that?!

OpenStudy (anonymous):

2x + 2y = 24.50 multiple by 2 = 4x + 4y = 49 3x + 4y = 37.25 multiple by 1 = 3x + 4y = 37.25 then 4x + 4y = 49 3x + 4y = 37.25 --------------- - x + 0 = 11.75 x = 11.75

OpenStudy (anonymous):

why did you multiply by 2 on the first one?

OpenStudy (anonymous):

to eliminate y because we juast need x

OpenStudy (anonymous):

ohhhh okay could you help me with another story problem?

OpenStudy (anonymous):

ok let me try it

OpenStudy (anonymous):

a movie is ocming to theathers and they are selling adult tickets for 10.50 and childrens tickers for 5.50. the theather sold 500 tickets opening night and made 4,700 dollars in sales. how many children attended the opening night?

OpenStudy (anonymous):

x = adult y = children x+y = 500 10.50x + 5.50y = 4700 because u need to find y, u have to eliminate x, can u do this?

OpenStudy (anonymous):

for the first equasion do i need to times everything by 10.50

OpenStudy (anonymous):

is the answer 50 kids?

OpenStudy (anonymous):

is the answer 50 kids?

OpenStudy (anonymous):

x+y = 500 | x 10.50| 10.50x + 10.50y = 5250 10.50x + 5.50y = 4700 |x1| 10.50x + 5.50y = 4700 -------------------- - 5y = 550 y = 110

OpenStudy (anonymous):

byt what about the first y?????

OpenStudy (anonymous):

y = children ticket

OpenStudy (anonymous):

hmm, are u got it?

OpenStudy (anonymous):

i dont get it at all...

OpenStudy (anonymous):

which part u dont understand?

OpenStudy (anonymous):

all of it.

OpenStudy (anonymous):

ok a movie is coming to theathers and they are selling adult tickets for 10.50 and childrens tickers for 5.50. the theather sold 500 tickets opening night and made 4,700 dollars in sales. how many children attended the opening night? 1st step to solve that kind of problem is choose a decision variable

OpenStudy (anonymous):

let x = # adult ticket and y = # children ticket then we have a constraint: 1. theatre capacity is 500 ticket so, x+y = 500 2. ticket prices for adult = 10.50 and children = 5.50 so , 10.50x + 5.50y = 4700 until this can u understand?

OpenStudy (anonymous):

yes, im jsut not understanding thennext part, like how you get the 2 equasions.

OpenStudy (anonymous):

yes, im jsut not understanding thennext part, like how you get the 2 equasions.

OpenStudy (anonymous):

yes, im jsut not understanding thennext part, like how you get the 2 equasions.

OpenStudy (anonymous):

theather capacity is 500 ticket so, we have to divide that theather for adult(x) and children (y) x+y = 500 ---> number of adult + number of children must be equal to 500

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

if i helped u can u give me medal or become my fans :D

OpenStudy (anonymous):

yeah i will!